# Babel-ImageNet: Massively Multilingual Evaluation of Vision-and-Language Representations

### v1

@@ 4 Comparative Evaluation of CLIP Models
Table 1: CLIP variants benchmarked on Babel-ImageNet with the following information: (i) the source (who trained the model), (ii) the text encoder, (iii) the image encoder (all variants of ViT (Dosovitskiy et al., 2021)), (iv) the training objective (CLIP: contrastive training as in Radford et al. (2021), LiT: locked image tuning Zhai et al. (2022), distill: MSE teacher distillation (Reimers and Gurevych, 2020)), (v) the amount of pretraining data (for “distill” the number of caption pairs, for CLIP/LiT the number of image-text pairs), and (vi) the number of languages covered by the training data.
Zero-Shot Image Classification: Setup. We adopt the ZS-IC setup of Radford et al. (2021): for an image-label pair, the image embedding is obtained directly from the image encoder; the label is inserted into 80 different prompt templates (from Radford et al. (2021)), each of which is independently embedded by the text encoder – the final label representation is then the mean of prompt embeddings. The class with the label embedding that is most similar to the image embedding (according to cosine similarity) is taken as the prediction; accuracy (top-1) is the evaluation metric.
Translating Prompts. We translate the 80 English prompts used by Radford et al. (2021) to our 92 languages using NLLB (Costa-jussà et al., 2022) (model: nllb-200-distilled-1.3B; see §C.2 for details). We show (see §D.1) that translated, language-specific prompts lead to better performance compared to using only the class labels or inserting them into the original English prompts. Moreover (see §D.2), we show that translated prompts yield similar performance as human-crafted prompts in ar, it, ja, and zh.
ZS-IC Results. Table 2 summarizes the results for the low-, mid-, and high-resource languages, alongside the English performance. Full per-language results for 92 languages can be found in the appendix (Table 7). Although differing subsets of ImageNet classes across language-specific benchmarks (see §3) prevent direct comparison of numbers between the language groups, these results make it abundantly clear that multilingual CLIP models perform dramatically worse (i) for high-resource languages than for English, and (ii) for low- and mid-resource languages than for high-resource languages. Note that this is despite the classification tasks a priori being easiest for low-resource languages (fewer than 333 classes) and hardest for English, where models must distinguish between all 1000 classes of ImageNet-1k.
Model low mid high en
OpenAI B-32 4.2 4.9 9.0 61.3
OpenCLIP XLMR B-32 15.0 31.0 39.7 62.8
M-CLIP XLMR-L B-32 25.7 32.8 33.3 42.6
M-CLIP XLMR-L B-16+ 25.8 34.5 36.0 46.4
M-CLIP mBERT B-32 14.8 19.3 18.9 29.2
ST mBERT B-32 9.2 15.1 17.1 38.2
M-CLIP XLMR-L L-14 28.1 37.7 39.5 51.6
AltCLIP XLMR-L L-14 14.2 21.1 33.6 69.9
OpenCLIP XLMR-L H-14 19.5 41.1 52.4 77.1
@@
Table 2: ZS-IC performance on Babel-ImageNet: average results for low-/mid-/high-resource languages and English. Bold: best result in each column, both between models with base (B) and large (L/H) image encoders.
@@
The OpenCLIP models, trained on the massive LAION5B data, yield good results for high- and mid-resource languages but perform poorly (in comparison with M-CLIP variants) for low-resource languages. In §D.4, we demonstrate that OpenCLIP performance strongly correlates with the distribution of languages in LAION5B: this would suggest that contrastive training (i.e., CLIP and LiT) leads to poor generalization across languages. In contrast, the best performance of M-CLIP models (with XLM-R as text encoder) on low-resource languages suggests that distillation-based training offers better cross-lingual generalization (yields best performance even for languages unseen in distillation training, see §D.3). We hypothesize that by aligning representations of captions in all other languages to the representations of corresponding English captions results in a more language-agnostic representation space. At the same time, in line with the “curse of multilinguality”, this improved generalization is paid with reduced quality of representations of high-resource languages, where M-CLIP models fall well behind OpenCLIP. This trade-off between cross-lingual generalization and per-language performance is best exemplified with AltCLIP: the model is exceptionally good for the 9 languages present in its large-scale distillation training (§D.3), yet performs (comparatively) poorly for most other languages – training on a very large dataset for only a few languages simply overwrites the XLM-R’s knowledge of other languages, obtained in its original pretraining.
The two mBERT-based models significantly underperform all other models. This is in part due to mBERT being generally a weaker multilingual text encoder than XLM-R Hu et al. (2020); Lauscher et al. (2020). On top of that, M-CLIP mBERT variants have been trained on less data than XLM-R-based counterparts (3M vs. 7M captions) and ST is distilled with parallel sentences that are not image captions.
@@ 5 Validating Babel-ImageNet
We perform two additional analyses that establish the validity of Babel-ImageNet as a benchmark: (1) how different number of classes affects performance and findings across languages and (2) how multilingual ZS-IC performance on Babel-ImageNet relates to multilingual image-text retrieval performance. We provide further analyses in the Appendix.
[…]
@@
Model low mid high en
OpenAI B-32 5.0 7.7 15.5 84.1
OpenCLIP XLMR B-32 16.4 39.5 56.2 85.3
M-CLIP XLMR-L B-32 26.1 41.2 47.3 66.3
M-CLIP XLMR-L B-16+ 26.6 42.2 48.6 70.7
M-CLIP mBERT B-32 15.9 26.5 30.3 49.9
ST mBERT B-32 10.4 22.3 28.6 60.1
M-CLIP XLMR-L L-14 28.2 45.2 51.9 72.2
AltCLIP XLMR-L L-14 15.7 27.5 45.7 90.9
OpenCLIP XLMR-L H-14 20.0 48.7 66.5 92.6
@@
Figure 3: Multilingual ZS-IC results when considering the same number of classes for each language (we report averages over 5 random subsets of 100 classes per language). For the 12 low-resource Babel-ImageNet languages with <100 classes, we take the single result on all classes.
Effect of number of classes on ZS-IC accuracy. Babel-ImageNet is an incomplete translation of the 1k ImageNet classes (see §3). Intuitively, classification tasks with fewer classes are easier and result in higher absolute performance for all models. We first analyze how the number of classes affects the ZS-IC performance on the English ImageNet-1K, for OpenAI CLIP models (B-32 and L-14). Figure 3 summarizes the results. The task difficulty (i.e., ZS-IC performance) is a log-linear function of the number of classes: this makes intuitive sense – moving from 50 to 100 classes increases the task difficulty much more than going from 900 to 950 classes.
We next fix the number of classes to 100 for all Babel-ImageNet languages (except for languages with <100 classes, for which we make no changes) and report the performance in Table 3 (for each language, we average the results over 5 different randomly selected subsets of 100 classes). [^9] It is not possible to select the exact same set of classes across all Babel-ImageNet languages because only one ImageNet-1k class (bee, Figure 1) has BabelNet translations in all 92 XLM-R languages. While in absolute terms the ZS-IC performance increases compared to full class sets (Table 2), and gaps between the language groups widen (especially between mid- and high-resource, and English to all three), our observations do not change: MCLIP (XLMR-L) still exhibits the best performance for low-resource languages, whereas OpenCLIP is the best for high-resource languages. This renders the (full) Babel-ImageNet a reliable benchmark for directly comparing multilingual VL models.
Multilingual ZS-IC vs. multilingual image-text retrieval. The aim of Babel-ImageNet, which measures multilingual ZS-IC performance, is to reflect the quality of the multilingual embedding spaces of VL models. Proving that this is the case implies showing that the models’ ZS-IC performance on Babel-ImageNet is indicative of their multilingual performance on tasks that – unlike ZS-IC itself – make sense in a multilingual formulation. The existing body of work commonly evaluates multilingual VL models in image-text retrieval. We thus compare how models’ performance on Babel-ImageNet correlates with their performance on three different multilingual image-text retrieval datasets: xFlickrCo (Bugliarello et al., 2022), XTD (Aggarwal and Kale, 2020), and XM3600 (Thapliyal et al., 2022), covering 7, 10, and 36101010For correlation analysis, we exclude fil, mi and quz as they are not within the 92 languages in Babel-ImageNet; we still report the models’ image-text retrieval performance for those languages in Table 9 languages, respectively. We use R@1 in text-to-image retrieval as the evaluation metric: it captures the percentage of examples where a correct image is top-ranked for a given caption. We report the full text-to-image retrieval results in the Appendix (Tables 8, 9, and 10).
@@
Figure 4: R@1 text-to-image retrieval results on three datasets plotted against Babel-ImageNet performance for 8 CLIP models (each dot denotes the performance of one model for one language) together with a linear regression estimate (95% CI).
@@
Figure 4 displays the retrieval results on xFlickrCo, XM3600, and XTD, respectively, against the ZS-IC accuracy on Babel-ImageNet: each dot represents one model-language combination. The plots reveal high correlation between the Babel-ImageNet and text-to-image retrieval scores across model-language pairs: $0.82$ for xFlickrCo, $0.87$ for XM3600, and $0.85$ for XTD. It is particularly positive that Babel-ImageNet shows the highest correlation with XM3600, with which it intersects in most languages (33). Balancing the number of classes to 100 for each language results in even higher correlations, e.g., $0.89$ against XM3600. These results confirm that Babel-ImageNet is a sensible benchmark for comparing proficiency of VL models for a multitude of languages not covered in task-specific (e.g., image-text retrieval) benchmarks.
@@ 6 Improving Multilingual CLIP for Low-Resource Languages
Finally, we improve the performance for low-resource languages by resorting to parameter-efficient fine-tuning with adapters (Houlsby et al., 2019; Pfeiffer et al., 2020b), trainable bottleneck layers that we insert into CLIP’s text encoder. In fine-tuning, we only update adapter parameters, keeping the original CLIP parameters frozen. We train a separate adapter for each (low-resource) language on top of the same CLIP model, which is computationally much cheaper than fine-tuning a specialized CLIP model for each language.
Setup. We train language-specific adapters on top of (a) OpenClip B-32 model (trained from scratch) and (b) M-CLIP XLMR-L B-32 (obtained via distillation). We experiment with three training objectives: English-target language distillation with (i) MSE and (ii) contrastive loss, and (iii) LiT on image-caption pairs. The former two require parallel data, whereas the latter requires images paired with target-language captions. For comparability between languages, we follow Carlsson et al. (2022) and sample 100K captions (with corresponding images) from the synthetic dataset provided by Li et al. (2022) and translate them automatically to all target languages with NLLB. We perform adapter-based specialization for 16 languages. One language adaptation run (i.e., one model-language-objective combination) took 3h on a single Nvidia RTX 3090 card (see §C.3 for details).
@@
Table 3: Results of adapter-based language adaptation of M-CLIP and OpenCLIP with three objectives (Text Contrastive, Text MSE, and LiT). Comparison against (i) the model without language adaptation and (ii) best-performing of all 8 CLIP models (see Table 1) for each language. Colors denote the size of change in performance w.r.t. original CLIP model: $\leq-5$, $\leq 0$, $\leq 5$, $\leq 10$, $\leq 20$, $>20$ (best viewed in color). We additionally report language statistics: the number of classes in Babel-ImageNet, the number of tokens used in XLM-R pre-training (in millions, log10), the number of examples in LAION5B (log10) and whether the language was used in M-CLIP training (True/False).
Results. Table 3 displays the results of language adaptation. For low-resource languages (xh, si, lo, my, and eu), we observe massive improvements, outperforming prior best results by wide margins. In contrast, the adaptation brings performance losses for high-resource languages (e.g., de and zh). We hypothesize that constraining the representation space of a target language to English representations is beneficial for low-resource languages with semantically poor initial representations, but detrimental for high-resource languages with semantically accurate initial representations. For both OpenCLIP and M-CLIP, adaptation with the MSE objective on parallel sentences yields the best results. Overall, the trends in performance changes from language adaptation are very similar between OpenCLIP and M-CLIP, despite the fact that they were obtained using very different training procedures and trained on datasets with different language distributions. This suggests that this commonality in language adaptation behavior stems from the initialization of the text encoder with XLM-R weights.

### v2

@@ 4 Benchmarking CLIP Models
Table 1: CLIP variants benchmarked with: (i) the source (who trained the model), (ii) the training objective (CLIP: contrastive training as in Radford et al. (2021), LiT: locked image tuning Zhai et al. (2022), distill: MSE teacher distillation (Reimers and Gurevych, 2020)), (iii) training data (for “distill” the number of caption pairs, for CLIP/LiT the number of image-text pairs), and (iv) the number of languages seen in training.
Model Param. v-low low mid high en
OpenAI B-32 151m 4.28 3.79 5.02 9.23 63.35
ST mBERT B-32 286m 6.23 9.72 15.33 17.44 39.45
M-CLIP mBERT B-32 330m 10.16 15.42 19.63 19.26 29.97
OpenCLIP XLMR-B B-32 366m 12.00 18.29 30.86 39.52 62.32
mSigLIP 371m 17.33 29.05 48.20 56.66 75.12
NLLB-SigLIP-base 507m 34.11 34.58 32.17 29.37 39.75
M-CLIP XLMR-L B-32 712m 18.52 26.40 33.47 34.11 44.06
M-CLIP XLMR-L B-16+ 769m 18.92 27.62 34.98 36.46 47.02
AltCLIP XLMR-L L-14 864m 12.67 16.98 21.32 33.97 73.36
M-CLIP XLMR-L L-14 988m 19.80 29.70 38.17 40.07 52.34
OpenCLIP XLMR-L H-14 1193m 13.77 23.57 41.03 52.23 76.95
NLLB-SigLIP-large 1195m 40.61 43.22 42.78 39.75 51.96
@@
Table 2: ZS-IC performance on Babel-ImageNet: average results for very-low-/low-/mid-/high-resource languages and English. Bold: best result in each column, both between models with base (B) and large (L/H) image encoders.
Models. We present the CLIP variants we benchmark on Babel-ImageNet (overview in Table 1). All use Vision Transformers (ViT) (Dosovitskiy et al., 2021) albeit of different sizes and with differently sized input patches (e.g., B-32 = Base Transformer with 32$\times$32-pixel patches).
OpenAI: The original CLIP (Radford et al., 2021), trained fully from scratch on 400M English image-caption pairs.
OpenCLIP: OpenCLIP (Ilharco et al., 2021) aims to replicate the OpenAI models using the public LAION datasets (Schuhmann et al., 2021, 2022). Two multilingual models have been trained: the B-32 model with the text encoder initialized with the weights of XLM-R-Base and the H-14 model initialized with XLM-R-Large. The B-32 variant is trained with the original contrastive CLIP objective, whereas the H-14 model was trained via locked image tuning (LiT, (Zhai et al., 2022)) in which the pretrained image encoder of the English H-14 OpenCLIP model is frozen and only the parameters of the text encoder are updated.
SentenceTransformer (ST): One of the first distillation-based multilingual CLIP-like models using the approach of Reimers and Gurevych (2020),666clip-ViT-B-32-multilingual-v1 with over 50M EN-X parallel sentences (X being one of 49 other languages) as supervision. They distill a (distilled) mBERT student (Devlin et al., 2019) from the English OpenAI B-32 teacher.
M-CLIP: Multilingual-CLIP (M-CLIP) Carlsson et al. (2022) is a model distilled using mBERT and OpenAI B-32 with translations of 3M English image captions to 69 languages as parallel supervision. Post-publication, they released a set of models777github.com/FreddeFrallan/Multilingual-CLIP with XLM-R-Large as the student and OpenAI B-32, L-14, and OpenCLIP B-16+ as teachers. These models were trained on 7M captions by Li et al. (2022), translated to 48 languages.
AltCLIP: This model by Chen et al. (2022b) distills an XLM-R-Large student with OpenAI L-14 as teacher, targeting 9 languages and using as training data a mix of machine-translated captions, multilingual captions sampled from LAION5B, and aligned English-X sentence pairs. After distillation, the authors additionally fine-tune the model via LiT using selected image-text pairs from LAION5B in the 9 target languages.
SigLIP: The multilingual mSigLIP Zhai et al. (2023) is a B-16-size CLIP model trained from scratch on Google’s multilingual WebLI dataset Chen et al. (2022a). They replace the softmax function used in the original CLIP with a sigmoid for more compute-efficient training.
NLLB-SigLIP: The NLLB-CLIP models Visheratin (2023) are a suite of CLIP models trained with LiT using 100k image-caption pairs translated fully for all 200 languages supported by NLLB Costa-jussà et al. (2022) for 20M examples in total. They use SigLIP B-16 and SO400M-14 as image encoders with the encoders of NLLB-600M-distilled and NLLB-1.3B-distilled as text encoders, which we call base and large, respectively.
Zero-Shot Image Classification Setup.
We adopt the ZS-IC setup of Radford et al. (2021): for an image-label pair, the image embedding is obtained directly from the image encoder; the label is inserted into 80 different prompt templates (from Radford et al. (2021)), each of which is independently embedded by the text encoder – the final label representation is then the mean of prompt embeddings. The class with the label embedding that is most similar to the image embedding (according to cosine similarity) is taken as the prediction; accuracy (top-1) is the evaluation metric.
Translating Prompts. We translate the 80 English prompts used by Radford et al. (2021) to our 100 languages using NLLB (Costa-jussà et al., 2022) (nllb-200-distilled-1.3B; see §B.2 for details). We show (see §C.1) that translated, language-specific prompts lead to better performance compared to using only the class labels or inserting them into the original English prompts. Moreover (see §C.2), we show that translated prompts yield similar performance as human-crafted prompts in ar, it, ja, and zh.
ZS-IC Results.
Table 2 summarizes the results for the four language groups, alongside the English performance. The full results for all 100 languages can be found in the Appendix (Table 8).
These results make it abundantly clear that multilingual CLIP models perform dramatically worse (i) for high-resource languages than for English, and (ii) for low- and mid-resource languages than for high-resource languages. Note that this is despite the classification tasks a priori being easiest for low-resource languages (under 333 classes) and hardest for English, where models must distinguish between all 1000 classes of ImageNet.
@@
The OpenCLIP models and mSigLIP, both trained on massive web-crawled corpora, yield good results for high- and mid-resource languages but perform poorly (in comparison to M-CLIP and NLLB-SigLIP) for low-resource languages. In §C.4, we demonstrate that OpenCLIP performance strongly correlates with the distribution of languages in LAION5B: this would suggest that contrastive training (i.e., CLIP and LiT) leads to poor cross-language generalization.
In contrast, the better performance of (XLM-R-based) M-CLIP models on low-resource languages suggests that distillation-based training offers stronger cross-lingual generalization (and yields best performance even for languages unseen in distillation training, see §C.3). We hypothesize that by aligning representations of captions in all other languages to the representations of corresponding English captions results in a more language-agnostic representation space. At the same time, in line with the “curse of multilinguality”, this improved generalization is paid with reduced quality of representations of high-resource languages, where M-CLIP models fall well behind.
The NLLB-SigLIP models, which combine a strong image encoder with the NLLB encoders trained on 200 languages, show impressive results for low- to mid-resource languages but underperform for the high-resource languages, suggesting again a trade-off between languages.
The trade-off between cross-lingual generalization and per-language performance is best exemplified with AltCLIP: the model is exceptionally good for the 9 languages present in its large-scale distillation training (§C.3), yet performs (comparatively) poorly for most other languages – training on a very large dataset for only a few languages simply overwrites the XLM-R’s knowledge of other languages, obtained in its original pretraining.
The two mBERT-based models significantly underperform all other models. This is in part due to mBERT being generally a weaker multilingual text encoder (Hu et al., 2020; Lauscher et al., 2020). On top of that, M-CLIP mBERT variants have been trained on less data than XLM-R-based counterparts (3M vs. 7M captions) and ST is distilled with parallel sentences that are not image captions.
@@ 5 Validating Babel-ImageNet
Model v-low low mid high en
ST mBERT B-32 6.23 11.83 24.36 32.65 64.79
M-CLIP mBERT B-32 10.16 18.30 29.86 34.49 54.21
OpenCLIP XLMR-B B-32 12.00 20.84 42.15 60.20 87.56
mSigLIP 17.33 32.20 59.83 75.83 93.58
NLLB-SigLIP-base 34.11 38.93 45.53 47.78 63.18
M-CLIP XLMR-L B-32 18.52 30.10 45.36 53.21 68.70
M-CLIP XLMR-L B-16+ 18.92 30.91 45.75 54.14 70.31
AltCLIP XLMR-L L-14 12.68 19.42 29.20 49.49 93.35
M-CLIP XLMR-L L-14 19.80 32.99 49.43 58.13 75.02
OpenCLIP XLMR-L H-14 13.77 26.32 52.65 71.65 94.09
NLLB-SigLIP-large 40.61 47.55 56.25 58.84 74.59
@@
Table 3: ZS-IC results with the same number of classes for each language (we report averages over 10 random subsets of 100 classes per language).
@@
Figure 3: R@1 text-to-image retrieval results on three datasets plotted against Babel-ImageNet performance (each dot denotes the performance of one model for one language) together with a linear regression estimate (95% CI).
We perform two additional analyses that establish the validity of Babel-ImageNet as a benchmark: (1) how different number of classes affects performance and findings across languages and (2) how ZS-IC performance on Babel-ImageNet relates to multilingual image-text retrieval performance. We provide further analyses in the Appendix.
Effect of number of classes on ZS-IC accuracy.
Babel-ImageNet is an incomplete translation of the ImageNet classes (see §3). Intuitively, classification with fewer classes is easier and results in higher absolute performance for all models. Our analysis of how the number of classes affects the ZS-IC performance on the English ImageNet for OpenAI CLIP models (B-32 and L-14) confirms this. Figure 2 shows that the task difficulty (i.e., ZS-IC performance) is log-linear in the number of classes: this makes intuitive sense – moving from 50 to 100 classes increases the difficulty more than going from 900 to 950 classes.
We next fix the number of classes to 100 for all Babel-ImageNet languages (except for languages with $<$100 classes, for which we make no changes) and report the performance in Table 3 (for each language, we average the results over 10 different randomly selected subsets of 100 classes). While in absolute terms the ZS-IC performance increases compared to full class sets (Table 2), and gaps between the language groups widen , our observations do not change: NLLB-SigLIP still exhibits the best performance for low-resource languages, whereas OpenCLIP and mSigLIP are still best for high-resource languages. This renders the (full) Babel-ImageNet a reliable benchmark for directly comparing multilingual VL models.
Multilingual ZS-IC & multilingual image-text retrieval.
The existing body of work commonly evaluates multilingual VL models in image-text retrieval. One goal of Babel-ImageNet, which measures multilingual ZS-IC performance, is to reflect (or, at least, give an estimate of) the quality of the multilingual embedding spaces of VL model. We thus compare how models’ performance on Babel-ImageNet correlates with their performance on three different multilingual image-text retrieval datasets: xFlickrCo (Bugliarello et al., 2022), XTD (Aggarwal and Kale, 2020), and XM3600 (Thapliyal et al., 2022), covering 7, 10, and 34888For correlation analysis, we exclude mi and quz as they are not under the 100 Babel-ImageNet languages. languages. We use R@1 in text-to-image retrieval as the evaluation metric: it captures the percentage of examples where a correct image is top-ranked for a given caption. We report the full retrieval results in the Appendix (Tables 10, 11, and 12).
@@
Figure 3 displays the retrieval results on xFlickrCo, XM3600, and XTD, respectively, against the ZS-IC accuracy on Babel-ImageNet: each dot represents one model-language combination. The plots reveal high correlation between the Babel-ImageNet and text-to-image retrieval scores across model-language pairs: $0.67$ for xFlickrCo, $0.75$ for XM3600, and $0.66$ for XTD. It is particularly positive that Babel-ImageNet shows the highest correlation with XM3600, with which it intersects in most languages (34). These results confirm that Babel-ImageNet is a sensible benchmark for comparing proficiency of VL models for a multitude of languages not covered by existing benchmarks.
The evaluation also reveals model-specific idiosyncrasies that consideration of either task alone would not have shown. For example, models that use a higher image resolution (M-CLIP B-16+, NLLB-SigLIPs) perform relatively better for retrieval, where the increased detail helps, than for ZS-IC. mSigLIP, potentially due to its small text encoder (compared to, e.g., XLM-R-large), is not as strong for retrieval as in ZS-IC. With the addition of ZS-IC to formerly retrieval-only multilingual evaluation, Babel-ImageNet allows for a more comprehensive evaluation of models.
@@ 6 Improving Multilingual CLIP for Low-Resource Languages
Table 4: Results of adapter-based language adaptation of M-CLIP and OpenCLIP with three objectives (Text Contrastive, Text MSE, and LiT). Comparison against the model before adaptation. Colors denote the size of change in performance w.r.t. original model: $\leq-5$, $\leq 0$, $\leq 5$, $\leq 10$, $\leq 20$, $>20$ (best viewed in color). We additionally report language statistics: the number of classes, the number of tokens in XLM-R pre-training (in millions, log10), the number of examples in LAION5B (log10) and whether the language was used in M-CLIP training (True/False).
Finally, we improve the performance for low-resource languages by resorting to parameter-efficient fine-tuning with adapters (Houlsby et al., 2019; Pfeiffer et al., 2020b), trainable bottleneck layers that we insert into the text encoder. We only update adapter parameters, keeping the original CLIP parameters frozen. We train a separate adapter for each language.
Setup.
We train language-specific adapters on top of (a) OpenClip B-32 model (trained from scratch) and (b) M-CLIP XLMR-L B-32 (obtained via distillation). We experiment with three training objectives: English-target language distillation with (i) MSE and (ii) contrastive loss, and (iii) LiT on image-caption pairs. The former two require parallel data, whereas the latter requires images paired with target-language captions. For comparability between languages, we sample 100K captions (with corresponding images) from the synthetic dataset provided by Li et al. (2022) and translate them automatically to all target languages with NLLB. We perform adapter-based specialization for 16 languages. One run (i.e., one model-language-objective combination) takes 3h on a single Nvidia RTX 3090 card (see §B.3 for details).
Results.
Table 4 displays the results. For low-resource languages (xh, si, lo, my, and eu), we observe massive improvements. In contrast, the adaptation brings performance losses for high-resource languages (e.g., de and zh). We hypothesize that constraining the representation space of a target language to English representations is beneficial for low-resource languages with semantically poor initial representations, but detrimental for high-resource languages with semantically accurate initial representations. For both OpenCLIP and M-CLIP, adaptation with the MSE objective on parallel sentences yields the best results. Overall, the trends in performance changes from language adaptation are very similar between OpenCLIP and M-CLIP, despite the fact that they were obtained using very different training procedures and trained on datasets with different language distributions. This suggests that this commonality in language adaptation behavior stems from the initialization of the text encoder with XLM-R weights.